OpenStudy (ashy98):

A.) Both are rational. B.) Both are irrational. C.)A is rational, but B is irrational. D.) A is irrational, but B is rational.

OpenStudy (ashy98):

Given the two expressions shown below: \[\sqrt{4}+\sqrt{25}\] \[\sqrt{4}+\sqrt{2}\] Which statement best describes the two expressions?

OpenStudy (ashy98):

@GracieBugg @rational @TylerD

OpenStudy (ashy98):

@PRAETORIAN.10

OpenStudy (rational):

\[\sqrt{4}+\sqrt{25} = 2 + 5 = 7\] Is \(7\) rational ?

OpenStudy (inkyvoyd):

is @rational being rational about 7?

OpenStudy (ashy98):

no?

OpenStudy (rational):

lol \[\large 7 = \dfrac{7}{1}\] now does it look rational ?

OpenStudy (rational):

rational number is a number which can be expressed as ratio of two integers

OpenStudy (rational):

All integers are rational because you can express them as ratio of two integers by simply putting \(1\) in the denominator

OpenStudy (ashy98):

so that means its c?

OpenStudy (rational):

Yes! A is rational and B is irrational

OpenStudy (ashy98):

yay thank you for helping!!!(:

OpenStudy (rational):

yw